April 26, 2017, Production & Analysis Working Group

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

The meeting will be on Wednesday April 26, 2017 at 1:30 pm EDT. For those people at Jefferson Lab, the meeting will be in room F326.

Meeting Connections

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/115815824.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 115815824.
  3. To join via phone, use one of the following numbers and the Conference ID: 115815824.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

Reconstruction & Analysis Studies Status


  1. Announcements
  2. Collaboration Meeting May 15-17, 2017
  3. Data Production
  4. Analysis Focus
  5. This Week's Studies
  6. Any other studies
  7. Upcoming Study Schedule

This Week's Topics

  • Talks: Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • Talks should be limited to 10 + 5 minutes.

Data Production

  1. Monitoring Update --- Thomas Britton
  2. Calibration Update --- Sean Dobbs
  3. Preparing for the last Spring 2016 production launch
  4. Preparing for the first Spring 2017 production launch
  5. Processing Update --- Alex Austregesilo

Analysis Focus: Total Cross Sections

  1. J/Ψ, φ --- Lubomir
  2. η, ρ, ω, φ update --- Simon
  3. φ update --- Thomas
  4. ω --- Christiano
  5. ρ yields for 2017 (mon_ver13) --- Alex
  6. Any others?

Studies: Beamline & Triggering

  1. Flux --- Justin
  2. Beam energy --- Beamline Group
  3. Polarization (TPOL & lineshape) --- Beamline Group
  4. Beam Asymmetries --- Alex Austregesilo
  5. Trigger emulation --- Alex Somov
  6. Triggering efficiency --- Alex Somov

Data Production Milestones

  1. Produce Spring 2016 "gold runs" as soon as alignment is ready
  2. Fri. May 5 - Finalize calibrations/software for low-rate running. Run final tests.
  3. Mon. May 8 -- Fri. May 12 (?) - Scicomp tape library down
  4. Fri. May 12 - Start production for low-rate running.
  5. Fri. May 19 - Finalize calibrations/software for high-rate running. Run final tests.
  6. Fri. May 26 - Start production for high-rate running.

Upcoming Study Schedule

  • Updates on experiment, simulation, and the comparison between the two.
  • Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • All talks should be limited to 10 + 5 minutes.

Next Week: Hit Efficiencies (i.e. Is the detector working?) Updates on experiment, simulation, and the comparison between the two.

  1. CDC Hit Efficiencies --- Naomi Jarvis
  2. FDC Hit Efficiencies --- Alex Austregesilo
  3. BCAL Hit Efficiencies --- Elton Smith
  4. FCAL Hit Efficiencies --- Jon Zarling

Week after next: Alignment & Track/Shower Efficiencies (i.e. Is the reconstruction working?) Updates on experiment, simulation, and the comparison between the two.

  1. Drift chamber alignment --- Mike Staib
  2. Tracking Efficiencies --- Paul Mattione
  3. BCAL Shower Efficiencies --- Elton Smith
  4. FCAL Shower Efficiencies --- Jon Zarling

The following week: Studies: Other Updates on experiment, simulation, and the comparison between the two.

  1. Track / Hit Matching: BCAL, FCAL, TOF, SC --- Paul Mattione, Simon
  2. Efficiencies: TOF, SC --- Beni, Mahmoud
  3. Means & Resolutions (time, energy, dE/dx): Tracking, BCAL, FCAL, SC, TOF
  4. Uncertainties: PID (BCAL, FCAL, TOF, dE/dx), Kinfit (BCAL, FCAL, tracking)
  5. Channel/Analysis Studies: Branching ratios, cross sections, SDMEs, beam asymmetries
  6. Other reconstruction/analysis issues


Data Production

  1. Sean reported that the timing alignment is essentially done, just cleaning up some outliers and understanding the timing of the high-rate runs. He is also looking at the systematics of these calibrations.
    • Mike S. has added his alignment updates and is working on updating the run-dependent time-to-distance relations.
    • Mark D. has an initial set of calibrations for the BCAL. Sean reported that he is obtaining much poorer resolutions with these calibrations. They are being investigated.
    • Sean pointed out one curiosity, that the TAGM base timing is much less stable than the TAGH. It should be checked to see if this is an artifact of the calibrations procedure or due to some real systematic effect that is being hidden by the calibration procedure.
  2. Alex A. reported that they have been running test jobs which produced good results, but crashed about 10% of the time. Once Paul M. has fixed a final bug today, a monitoring launch on the 2016 data will be performed. The schedule for the next 2017 data monitoring launch is being discussed.

Analysis Focus: Total Cross Sections

  1. Lubomir has been looking at a combination of the available 2016 and 2017 data, and is able to perform the analysis in bins of E(gamma). He will show some more detailed results at the dilepton meeting tomorrow.
  2. Simon showed some updated results for cross sections of phi and omega decays into 3pi and 3gamma. He has improved the robustness of the fits, although it appears that for the omega -> pi0 gamma decay, signal events are being absorbed into the background. The huge bin-to-bin variations he showed previously are gone. The phi cross sections are more-or-less consistent with expectations within statistical uncertainty. The omega cross sections have several features that need to be understood.
    • Paul M. pointed out that these results have a cut on the kinematic fit confidence level and that could be affecting the results. Eugene pointed out that the backgrounds are small enough that a cut-and-count approach could be used as well.
    • Thomas showed the first results for his phi -> K+K- cross section. The results are very preliminary, but are in the right ballpark and are ~30-50% larger than Simon's results. Understanding the simulation results and error calculations are the next steps.
    • Alex A. showed a summary of the rho rates from 2017 data using the latest (ver13) monitoring launch. The results are more or less stable for the low-rate running, and vary more for the high-rate running, where the calibrations were not as mature at the time of the launch. He will look more closely at the systematics of these measurements and the asymmetry results from them.

Studies: Beamline & Triggering

  1. Justin is finalizing flux numbers for 2016 to be stored in CCDB and will provide tools for working with this information.
  2. The current endpoint beam energy values in the CCDB are considered to be good to within 30 MeV. The goal is reduce this uncertainty to 10 MeV. Alexandre Deur will make a presentation at the Collaboration meeting.
  3. Nathan will be generating the ROOT trees needed for the full TPOL analysis of Spring 2017 data based on the PS skims and the fully aligned detectors.
  4. Alex A. will give an update on beam asymmetries for Spring 2017 likely next week.
  5. Alex S. is making progress on understanding the trigger efficiency using production data instead of raw data, and implementing it in the code.